The odometer on the family car read 15951 when the driver noticed that the number was palindromic. "Curious," said the driver to herself. "It will be a long time before that happens again." But an hour and 50 minutes later, the odometer showed a new palindromic number. How fast was she traveling?

Answer by ptaylor(1661)

The next palindromic number would be: 16061
Miles travelled to reach next palindromic number: 16061-15951=110 mi
distance(d)=rate(r) times time(t) or d=rt; t=d/r and r=d/t
110=r(1 5/6hr)=r(11/6) multiply both sides by 6
660=11r divide both sides by 11
r= 60mph------------------------------her speed
Hope this helps ----ptaylor